April took out a $10000 loan to buy a car. The loan has an interest rate of 2.1% for 4 years with monthly payments. What will April's monthly payments be? (Round up to the cents place.)
With one year left on the loan, what is April's remaining balance? (Round up to the cents place.)

Answer by math_tutor2020(3817)


L = 10,000 = loan amount
i = interest rate per month = 0.021/12 = 0.00175
n = number of months = 12*4 = 48

P = monthly payment
P = (L*i)/( 1-(1+i)^(-n) )
P = (10,000*0.00175)/( 1-(1+0.00175)^(-48) )
P = 217.387950449613
P = 217.39

m = number of months left = 12
B = remaining balance with m months left
B = P*(1-(1+i)^(-m))/i
B = 217.387950449613*(1-(1+0.00175)^(-12))/(0.00175)
B = 2579.22270192667
B = 2579.22

The monthly payment and remaining balance can be confirmed when using a calculator such as this one
https://www.calculator.net/loan-calculator.html
After you type in the proper values, and hit "calculate", it should show the monthly payment being $217.39
Also, click on "View Amortization Table". Note the balance at the end of month 36 (and at the beginning of month 37) is exactly $2,579.22

-----------------------------------------------------------

Answers:
monthly payment = $217.39
Remaining balance with one year left = $2,579.22